# Challenges Faced by Large Language Models in Solving Multi-Agent Flocking

### v1

@@ Abstract
Flocking is a behavior where multiple agents in a system attempt to stay close to each other while avoiding collision and maintaining a desired formation. This is observed in the natural world and has applications in robotics, including natural disaster search and rescue, wild animal tracking, and perimeter surveillance and patrol. Recently, large language models (LLMs) have displayed an impressive ability to solve various collaboration tasks as individual decision-makers. Solving multi-agent flocking with LLMs would demonstrate their usefulness in situations requiring spatial and decentralized decision-making. Yet, when LLM-powered agents are tasked with implementing multi-agent flocking, they fall short of the desired behavior. After extensive testing, we find that agents with LLMs as individual decision-makers typically opt to converge on the average of their initial positions or diverge from each other. After breaking the problem down, we discover that LLMs cannot understand maintaining a shape or keeping a distance in a meaningful way. Solving multi-agent flocking with LLMs would enhance their ability to understand collaborative spatial reasoning and lay a foundation for addressing more complex multi-agent tasks. This paper discusses the challenges LLMs face in multi-agent flocking and suggests areas for future improvement and research.
@@ 2.1 Multi-agent Flocking
Multi-agent flocking refers to the coordinated movement of a group of independent agents in a way that resembles the natural flocking behavior of birds. This topic is studied to understand how groups of agents can work together without a leader to achieve the desired formation. The paper [19] highlights the three flocking rules of Reynolds namely: Flock Centering; Collision Avoidance and Velocity Matching. Based on these rules, Olfati-Saber has demonstrated three different flocking algorithms [15] for $\alpha$ agents. An $\alpha$ agent’s main objective is to create an $\alpha$-lattice with its neighboring $\alpha$-agents. The first Algorithm generates spatial order in flocks, incorporating all three rules of Reynolds in one equation, leading to regular fragmentation. The second Algorithm leads to flocking in free space, with an added consideration for $\gamma$ agents for group objectives. The third Algorithm combines the first two algorithms and obstacle avoidance capabilities provided by virtual agents known as $\beta$ agents. Recent work in multi-agent flocking includes studying the two collective behaviors like swarming and flocking [20]. Flocking involves agents aligning their motion, whereas swarming involves them gathering together to coordinate their spatial arrangement. It is observed that inter-particle avoidance is a necessary condition for flocking. Li et al. [11] proposed the artificial potential-based approach for the migration and trajectory tracking of multi-agent systems. Zhu et al. [33] utilized deep reinforcement learning to learn multi-agent flocking control in complex environments with dynamic obstacles. Multi-robot flocking control using Multi-Agent Twin Delayed Deep Deterministic Policy Gradient (MATD3) has been studied to solve the issue of overestimation bias [30]. Chen et al. [26] has discussed how dividing the flocking group into subgroups can help to address the problem of maintaining social distancing among them and allow for better coordination. This study [12] proposes a discrete-time flocking control model for multi-robot systems with random link failures, ensuring practical implementation and achieving flocking in expectation with improved stability conditions.
@@ 3.1.1 Multi-agent flocking
Multi-agent flocking is a well-studied problem. Reynolds[19] proposed the first approach to simulate the flocking behavior of multiple agents for computer animation purposes in 1986. Later, an algorithm for mobile agents with a double integrator dynamic model was proposed by Tanner et al. [24]. Then, Olfati-Saber proposed a decentralized approach to the multi-agent flocking problem with similar double integrator dynamics [15]. Olfati-Saber formulated the control laws consist a combination of attractive/repulsive, alignment, and leading forces, which tackled the collision avoidance and cohesion of the group, the common heading direction for an aggregate motion, as well as the potential goal location for the flock to reach.
@@ 3.2 LLMs for flocking
Here, the “agent_role” is the role of each agent in flocking. The “game_description” explains the rules and goals of the game. The game aims to form a flock with all agents by keeping the Boid flocking rules in mind. The hyperparameters specified are the positions of the agent and its neighboring agents, the flock shape, the safe distance, and the maximum velocity (maximum units the agent can move). The agent role and game description are provided to the LLM in the first round. The “round_description”, provided in the subsequent rounds, contains the updated positions and reminds the LLM to keep maximum velocity in mind. The “output_format” is ‘Reasoning: …, Position: [x,y]’ in which the LLM output is expected to be structured. Note that, here we only provide an upper bound for the velocity but do not directly control it. The reason is twofold. Firstly, adding velocity control along with position control gives another layer of complexity for the LLMs to reason and understand, making it difficult for them to interpret the decision-making process. Secondly, we want to provide minimal information to the LLMs to rely more on them to reason and make decisions.
@@ 4 Cause of Failures
We first test multi-agent flocking with a group size of five to form a specific pattern. Each agent starts from a random position within the world and has its own LLM to decide the motion. The prompt contains the current positions of the agent and its neighbors, as well as the flock pattern the multi-agent group needs to form and the safe distance each agent must maintain. The LLM will make decisions based on the limited information provided and generate the new position the agent should go to. We specify different flock patterns, including circle, $\alpha$-lattice as described in [15], and V-shape. We perform ten tests on each flock pattern. We use mean absolute error (MAE) to measure the performance of the test. The mean absolute error is calculated by
@@
Three-agent Tests. After the failed attempt to form a flock with five agents, we set up a simpler scenario, where three agents are staying at the desired distance between neighbors. The desired distance is five units between any two agents. Our expectation for this scenario is the three agents will stay close together and form a triangle with an equal distance between each agent. The quantitative result with ten tests is shown in Fig. 1(b), showing the MAE does not decrease to the desired value. The trajectory of the agents is similar to the flock of five agents, in which agents gather together instead of maintaining the desired distance. Different from the five-agent flocking scenario, flocking with three agents does not require the LLM to process complicated shape information and coordination. The triangle pattern is naturally formed when an agent tries to maintain a desired distance from the other two neighbor agents. We then question if the LLM is capable of reasoning the distance and making corresponding decisions, which is crucial for flock formation.
Two-agent Tests. To test the ability to reason and understand distance, we set up the tests with two agents starting from random locations and trying to keep a desired distance between them. The MAE for the ten tests is shown in Fig. 1(c). For all the tests, the desired distance is 10 units between two agents. Unlike the previous tests, where the agents gather to form a consensus, the two agents keeping a certain distance scenario is more random and spontaneous. Some tests show the consensus behavior, but other tests show the agents moving away from each other instead of staying at a certain distance. To further simplify the problem, we make one of the agents stationary, which means only one agent is moving around to keep the desired distance. The MAE of ten tests is shown in Fig. 1(d). A similar behavior is observed here in that the moving (or active) agent. Instead of keeping the desired distance, drives further away from the stationary agent.

### v2

# Corresponding author

@@ Abstract
Flocking is a behavior where multiple agents in a system attempt to stay close to each other while avoiding collision and maintaining a desired formation. This is observed in the natural world and has applications in robotics, including search and rescue, wild animal tracking, and perimeter surveillance. Recently, large language models (LLMs) have displayed an impressive ability to solve various collaboration tasks as individual decision-makers. Solving multi-agent flocking with LLMs would demonstrate their usefulness in situations requiring spatial and decentralized decision-making. Yet, when LLM-powered agents are tasked with implementing multi-agent flocking, they fall short of the desired behavior. After extensive testing, we find that agents with LLMs as individual decision-makers typically opt to converge on the average of their initial positions or diverge from each other. After breaking the problem down, we discover that LLMs cannot understand maintaining a shape or keeping a distance in a meaningful way. Solving multi-agent flocking with LLMs would enhance their ability to understand collaborative spatial reasoning and lay a foundation for addressing more complex multi-agent tasks. This paper discusses the challenges LLMs face in multi-agent flocking and suggests areas for future improvement and research.
@@ 2.1 Multi-agent Flocking
Multi-agent flocking refers to the coordinated movement of a group of independent agents in a way that resembles the natural flocking behavior of birds. This topic is studied to understand how groups of agents can work together without a leader to achieve the desired formation. The paper [19] highlights the three flocking rules of Reynolds namely: Flock Centering, Collision Avoidance, and Velocity Matching. Based on these rules, Olfati-Saber has demonstrated three different flocking algorithms [15] for $\alpha$ agents. An $\alpha$ agent’s main objective is to create an $\alpha$-lattice with its neighboring $\alpha$-agents. The first Algorithm generates spatial order in flocks, incorporating all three rules of Reynolds in one equation, leading to regular fragmentation. The second Algorithm leads to flocking in free space, with an added consideration for $\gamma$ agents for group objectives. The third Algorithm combines the first two algorithms and obstacle avoidance capabilities provided by virtual agents known as $\beta$ agents. Recent work in multi-agent flocking includes studying the two collective behaviors like swarming and flocking [20]. Flocking involves agents aligning their motion, whereas swarming involves them gathering together to coordinate their spatial arrangement. It is observed that inter-particle avoidance is a necessary condition for flocking. Li et al. [11] proposed the artificial potential-based approach for the migration and trajectory tracking of multi-agent systems. Zhu et al. [33] utilized deep reinforcement learning to learn multi-agent flocking control in complex environments with dynamic obstacles. Multi-robot flocking control using Multi-Agent Twin Delayed Deep Deterministic Policy Gradient (MATD3) has been studied to solve the issue of overestimation bias [30]. Chen et al. [26] has discussed how dividing the flocking group into subgroups can help to address the problem of maintaining social distancing among them and allow for better coordination. This study [12] proposes a discrete-time flocking control model for multi-robot systems with random link failures, ensuring practical implementation and achieving flocking in expectation with improved stability conditions.
@@ 3.1.1 Multi-agent flocking
Olfati-Saber formulated the control laws consist a combination of attractive/repulsive, alignment, and leading forces [15], which tackled the collision avoidance and cohesion of the group, the common heading direction for an aggregate motion, as well as the potential goal location for the flock to reach.
@@ 3.2 LLMs for flocking
Here, the “agent_role” is the role of each agent in flocking. The “game_description” explains the rules and goals of the game. The game aims to form a flock with all agents by keeping the Boid flocking rules in mind. The hyperparameters specified are the positions of the agent and its neighboring agents, the flock shape, the safe distance, and the maximum velocity (maximum units the agent can move). The agent role and game description are provided to the LLM in the first round. The “round_description”, provided in the subsequent rounds, contains the updated positions and reminds the LLM of the maximum velocity. The “output_format”, ‘Reasoning: …, Position: [x,y]’, describes the expected output structure. Note that, here we only provide an upper bound for the velocity but do not directly control it. The reason is twofold. Firstly, adding velocity control along with position control gives another layer of complexity for the LLMs to reason and understand, making it difficult for them to interpret the decision-making process. Secondly, we want to provide minimal information to the LLMs to rely more on them to reason and make decisions.
@@ 4 Cause of Failures
Three-agent Tests. After the failed attempt to form a flock with five agents, we set up a simpler scenario, where three agents are staying at the 5 units’ desired distance between neighbors. Our expectation for this scenario is the three agents will stay close together and form a triangle with an equal distance between each agent. The quantitative result with ten tests is shown in Fig. 1(b), showing the MAE does not decrease to the desired value. The trajectory of the agents is similar to the flock of five agents, in which agents gather together instead of maintaining the desired distance. Different from the five-agent flocking scenario, flocking with three agents does not require the LLM to process complicated shape information and coordination. The triangle pattern is naturally formed when an agent maintains a desired distance from the other two neighbor agents. We then question if the LLM is capable of reasoning the distance and making corresponding decisions, which is crucial for flock formation.
Two-agent Tests. To test the ability to reason and understand distance, we set up the tests with two agents starting from random locations and trying to keep a desired distance between them. The MAE for the ten tests is shown in Fig. 1(c). For all the tests, the desired distance is 10 units between two agents. Unlike the previous tests, where the agents gather to form a consensus, the two agents keeping a certain distance scenario is more random and spontaneous. Some tests show the consensus behavior, but other tests show the agents moving away from each other instead of staying at a certain distance. To further simplify the problem, we make one of the agents stationary, which means only one agent is moving to keep the desired distance. The MAE of ten tests is shown in Fig. 1(d). A similar behavior is observed here in that the moving (or active) agent. Instead of keeping the desired distance, drives further away from the stationary agent.